OpenStudy (mwhit2101):

8TH GRADE ALGEBRA Please help fast! Medal + Fan for best answer. Please help me. I must give the equations of two lines that are parallel to y=1/2x+5

OpenStudy (mwhit2101):

\[y=\frac{ 1 }{ 2 }x+5\]

OpenStudy (anonymous):

Parallel lines have same slope.

OpenStudy (anonymous):

First, find the slope of given line..

OpenStudy (mwhit2101):

m = 1/2

OpenStudy (mwhit2101):

Yes?

OpenStudy (anonymous):

Good..

OpenStudy (anonymous):

So, the other two lines must have same slope, right?

OpenStudy (mwhit2101):

Yes

OpenStudy (anonymous):

They must be like: \[y = \frac{1}{2}x + c_1\] \(y = \frac{1}{2}x + c_2\)

OpenStudy (mwhit2101):

Oh um what is c1 and c2?

OpenStudy (mwhit2101):

I havent done those I don't think

OpenStudy (anonymous):

Like in your original equation, you have \(5\), right?

OpenStudy (anonymous):

\(5\) in that equation is your c, \(c\) is called \(y\)-intercept..

OpenStudy (mwhit2101):

Oh, okay, got it. It's always been b for y-intercept but c works too

OpenStudy (mwhit2101):

What next?

OpenStudy (anonymous):

In INDIA, it is \(c\).. :P

OpenStudy (mwhit2101):

Oh ok well that's fine

OpenStudy (anonymous):

So, \(c_1\) and \(c_2\), you can choose yourself..

OpenStudy (mwhit2101):

Choose what, a number?

OpenStudy (anonymous):

They will just shift your original line upward or downward.. Yes a number..

OpenStudy (mwhit2101):

8

OpenStudy (anonymous):

Your original equation is: